Resistance of a metal wire of lenght 2m is 26 ohm at 20°c if the diameter of the wire is 0.3 mm ,what will be the resistivity of metal at that temperature?

Grade:10

1 Answers

Shreya D Nanda
6 years ago
HI,
We can use the formula :R= \rho L/A , where, R is resistance of the wire, L its length, A its cross-sectional area and \rho [rho] is the resistivity of the wire.
 From the data given, R=26 \OmegaL=2m  and    A=\pi[0.3*10^{-3}]^{2} m^{}2    i.e, A= 9*10^{-8}\pi m^{2}
 
\rho =RA/L  \Rightarrow \rho = 26 * 9*10^{-8}\pi /2
\therefore \rho = 367.38 * 10^{-8} \Omega m
or \rho = 3.674 * 10^{-6} \Omega m
 
Hope it helps, thanks..